7)Find and compare the future value after two years of a deposit of $100 attracting interest at a rate a) annually and b) semiannually.
8)Which will deliver a higher future value after one year, a deposit of $1, 000 attracting interest at 15% compounded daily, or at 15.5% compounded semi-annually?
9)What initial investment subject to annual compounding at 12% is needed to produce $1, 000 after two years?
10)Which will deliver a higher future value after one year, a deposit of $1, 000 attracting interest at 15% compounded daily, or at 15.5% compounded semi-annually?
11)What initial investment subject to annual compounding at 12% is needed to produce $1, 000 after two years?
12)Find the present value of $100, 000 to be received after 100 years if the interest rate is assumed to be 5% throughout the whole period and a) daily or b) annual compounding applies.
1
Expert's answer
2020-06-09T18:59:43-0400

7) Assuming an interest rate of 15%


compounding interest formula is as follows;

"FV= PV* (1+r\/n){^n}{^t}"

FV is future value

PV is present value

r is interest

n is number of compounding in a year

t is the total number of years


a)compounded annually


"FV= 100 * (1+0.15\/1)^{1*2}"

"FV=132.25\/"


b) compounded semi-annually


"FV= 100 * (1+0.15\/2)^{2*2}"

"FV=133.55\/"


8)

a)compounded daily at 15%


"FV= 1,000 * (1+0.15\/365)^{365*1}"

"FV=1,161.80"


b) compounded semi-annually at 15.5%

"FV= 1,000 * (1+0.155\/2)^{2*1}"

"FV=1,161.01\/"

Compounding daily is slightly higher


9)

"1,000= PV* (1+0.12\/1)^{2*1}"

"PV=797.19"


10) Is a repetition of question 8

11) Is a repetition of question 9


12)

a)compounded daily


"100,000= PV* (1+0.05\/365)^{100*365}"

"PV=674.03\/"


b)compounded daily


"100,000= PV* (1+0.05\/1)^{100*1}"

"PV=760.45\/"